A coin was tossed five times. whats the probability that it will show heads or tails atleast once.
Answer by stanbon(75887)

A coin was tossed five times. whats the probability that it will show heads or tails at least once.
---
P(at least one head) = 1 - P(no heads) = 1-(1/2)^5 = 31/32
=============================
Cheers,
Stan H.
============